# The Fast and the Frame-Dragging: Efficient waveforms for asymmetric-mass eccentric equatorial inspirals into rapidly-spinning black holes

### v1

@@ Appendix F Marginal posterior distributions for science-case sources
In this Appendix, we show corner plots of the marginal posterior distributions obtained from the parameter estimation analyses performed in Section V.6. Figures 28, 29 and 31 correspond to the analysis of EMRIs with parameters given by rows 1, 2 and 5 of Table 1 respectively. We do not observe any significant biases in the intrinsic parameters of these sources for $\kappa=10^{-2}$, which is expected (based on Eq. 15) given that these sources have SNRs $\leq 50$.
@@
Similarly, Figs. 32 and 30 correspond to the analysis of IMRIs with parameters given by rows 3 and 4 of Table 1 respectively. As the SNRs of these sources is higher (500 and 300 respectively), analyses with $\kappa=10^{-2}$ failed to converge due to large mismodelling errors shifting the posterior bulk far from the parameters of the injections. For $\kappa=10^{-3}$, biases in intrinsic parameters are generally below $1\,\sigma$, but more significant biases are observed in the recovery of extrinsic parameters.
@@
Figure 28: Posterior distributions of the intrinsic parameters for an EMRI with parameters given in the first row of Table 1 for two values of the mode selection threshold. The blue (orange) posteriors refer to a threshold set to $\kappa=10^{-5}$ ($10^{-2}$). 2D (1D) shaded areas refer to 1- and 2-sigma (95%) credible regions. For the intrinsic parameters, we plot the quantities $\Delta X=X-X_{\rm true}$ for visual clarity.
@@
Figure 29: Posterior distributions of the intrinsic parameters for an EMRI with parameters given in the second row of Table 1 for two values of the mode selection threshold. The blue (orange) posteriors refer to a threshold set to $\kappa=10^{-5}$ ($10^{-2}$). 2D (1D) shaded areas refer to 1- and 2-sigma (95%) credible regions. For the intrinsic parameters, we plot the quantities $\Delta X=X-X_{\rm true}$ for visual clarity.
@@
Figure 30: Posterior distributions of the intrinsic parameters for an IMRIs with parameters given in the fourth row of Table 1 for two values of the mode selection threshold. The blue (orange) posteriors refer to a threshold set to $\kappa=10^{-5}$ ($10^{-3}$). 2D (1D) shaded areas refer to 1- and 2-sigma (95%) credible regions. For the intrinsic parameters, we plot the quantities $\Delta X=X-X_{\rm true}$ for visual clarity. We also include the luminosity distance $d_{\mathrm{L}}$ to show the bias induced by the mode selection threshold.
@@
Figure 31: Posterior distributions of the intrinsic parameters for an EMRI with parameters given in the fifth row of Table 1 for two values of the mode selection threshold. The blue (orange) posteriors refer to a threshold set to $\kappa=10^{-5}$ ($10^{-2}$). 2D (1D) shaded areas refer to 1- and 2-sigma (95%) credible regions. For the intrinsic parameters, we plot the quantities $\Delta X=X-X_{\rm true}$ for visual clarity.
@@
Figure 32: Posterior distributions for an IMRI with parameters given in the third row of Table 1 for two values of the mode selection threshold. The blue (orange) posteriors refer to a threshold set to $\kappa=10^{-5}$ ($10^{-3}$). 2D (1D) shaded areas refer to 1- and 2-sigma (95%) credible regions. For the intrinsic parameters, we plot the quantities $\Delta X=X-X_{\rm true}$ for visual clarity. The $\kappa=10^{-3}$ posteriors yield significant biases in the extrinsic parameters due to the high SNR of this source (500).

### v2

@@ Appendix F Marginal posterior distributions for science-case sources
In this Appendix, we show corner plots of the marginal posterior distributions obtained from the parameter estimation analyses performed in Section˜V.6. Figures˜28, 29 and 31 correspond to the analysis of EMRIs with parameters given by rows 1, 2 and 5 of Table˜1 respectively. We do not observe any significant biases in the intrinsic parameters of these sources for $\kappa=10^{-2}$, which is expected (based on Eq.˜15) given that these sources have SNRs $\leq 50$.
@@
Similarly, Figs.˜32 and 30 correspond to the analysis of IMRIs with parameters given by rows 3 and 4 of Table˜1 respectively. As the SNRs of these sources is higher (500 and 300 respectively), analyses with $\kappa=10^{-2}$ failed to converge due to large mismodelling errors shifting the posterior bulk far from the parameters of the injections. For $\kappa=10^{-3}$, biases in intrinsic parameters are generally below $1\,\sigma$, but more significant biases are observed in the recovery of extrinsic parameters.
@@
Figure 28: Posterior distributions of the intrinsic parameters for an EMRI with parameters given in the first row of Table˜1 for two values of the mode selection threshold. The blue (orange) posteriors refer to a threshold set to $\kappa=10^{-5}$ ($10^{-2}$). 2D (1D) shaded areas refer to 1- and 2-sigma (95%) credible regions. For the intrinsic parameters, we plot the quantities $\Delta X=X-X_{\rm true}$ for visual clarity.
@@
Figure 29: Posterior distributions of the intrinsic parameters for an EMRI with parameters given in the second row of Table˜1 for two values of the mode selection threshold. The blue (orange) posteriors refer to a threshold set to $\kappa=10^{-5}$ ($10^{-2}$). 2D (1D) shaded areas refer to 1- and 2-sigma (95%) credible regions. For the intrinsic parameters, we plot the quantities $\Delta X=X-X_{\rm true}$ for visual clarity.
@@
Figure 30: Posterior distributions of the intrinsic parameters for an IMRIs with parameters given in the fourth row of Table˜1 for two values of the mode selection threshold. The blue (orange) posteriors refer to a threshold set to $\kappa=10^{-5}$ ($10^{-3}$). 2D (1D) shaded areas refer to 1- and 2-sigma (95%) credible regions. For the intrinsic parameters, we plot the quantities $\Delta X=X-X_{\rm true}$ for visual clarity. We also include the luminosity distance $d_{\mathrm{L}}$ to show the bias induced by the mode selection threshold.
@@
Figure 31: Posterior distributions of the intrinsic parameters for an EMRI with parameters given in the fifth row of Table˜1 for two values of the mode selection threshold. The blue (orange) posteriors refer to a threshold set to $\kappa=10^{-5}$ ($10^{-2}$). 2D (1D) shaded areas refer to 1- and 2-sigma (95%) credible regions. For the intrinsic parameters, we plot the quantities $\Delta X=X-X_{\rm true}$ for visual clarity.
@@
Figure 32: Posterior distributions for an IMRI with parameters given in the third row of Table˜1 for two values of the mode selection threshold. The blue (orange) posteriors refer to a threshold set to $\kappa=10^{-5}$ ($10^{-3}$). 2D (1D) shaded areas refer to 1- and 2-sigma (95%) credible regions. For the intrinsic parameters, we plot the quantities $\Delta X=X-X_{\rm true}$ for visual clarity. The $\kappa=10^{-3}$ posteriors yield significant biases in the extrinsic parameters due to the high SNR of this source (500).